OpenStudy (anonymous):

What is the y-intercept in this linear equation?y = -4x + 4

OpenStudy (anonymous):

The y intercept is defined where x = 0. What does that give us? ...

OpenStudy (luigi0210):

set x equal to zero

OpenStudy (anonymous):

y=-4(0)+4 y=0+4 y=4

OpenStudy (luigi0210):

There you go!

OpenStudy (anonymous):

well looky there

OpenStudy (johnweldon1993):

Also good to note that with a linear equation like this, you can use the slope intercept equation y = mx + b and 'b' is your y-intercept so your y = -4x + 4 y = mx + b as you can see (and have shown above...the y-intercept is 4